# A new indeterminacy-based quantum theory

### v4

@@ 2 Main features: four conditions and two hypotheses
One might also object that the inflationary paradigm has its problems, which places EnDQT in problematic foundations. [^44] See, e.g., Ijjas et al., (2013), Dawid & McCoy (2023) and references therein. However, note that the SDCs-starting-hypothesis is a placeholder for current and future physics. Furthermore, whatever theory substitutes inflation, it should deal with the problems it pertains to solve. To solve those problems, it’s plausible to consider that some features shared with the initiators will likely arise. This is because it’s plausible to expect that in a future theory, it’s likely that we will also have a rapid expansion of the universe in the early stages that doesn’t occur anymore. Such rapid expansion is likely due to some set of entities E with a considerable influence that don’t manifest themselves anymore or so much at least. Furthermore, the SDCs-starting hypothesis can also be supplemented with another hypothesis if it turns out that there are phenomena that need to be explained via widespread initiators. So, EnDQT, conceived more broadly, is a view whose truth doesn’t just rely on inflation. [^45] It’s conceivable that there are other alternative mechanisms beyond inflation, see Appendix C.
@@
I have mentioned above some predictions that EnDQT provides, such as the dissolution of SDCs and those that are a consequence of the SDCs-inflationary hypothesis. Another prediction is the following: as we have seen with the example above, adopting the conservative determination conditions (CDCs) generates constraints on how SDCs are formed and new predictions. Decoherence timescales roughly serve as an indicator for the timescale it takes for environments of a system to decohere that system on average, where that system ends up having specific determinate values (that are observed in the lab). Given the SDCs-decoherence hypothesis, the CDCs predict that the decoherence timescale that we empirically observe of a kind of system $Z$ by a kind of system $Y$ should be superior or of the same order as the decoherence timescale of $Y$ by a kind of system $X$, where $Y$ is typically decohered by $X$ before $Y$ decoheres $Z$, and where the interaction between $X$ and $Y$ starts first. Otherwise, contrary to what is assumed by the CDCs, we can have situations where $Z$ will have a determinate value first (due to $Y$), then $Y$ will have a determinate value due to $X$. Since the decoherence timescales are typically empirically determined, a further analysis of the current empirically determined decoherence timescales is needed to see if they agree with the predictions of the CDCs.
@@
The predictions of this condition are supported in the case that $Y$ is a macrosystem (e.g., measurement devices), and $Z$ is a microsystem. This is because macroscopic systems have decoherence timescales much shorter than the microscopic systems that they can decohere. [^46] The cross section for larger systems is larger than the one for smaller systems. Moreover, the decoherence rate of a quantum system, which is the inverse of the decoherence timescale, is proportional to their cross-section, as well as the flux of systems of the environment. See the collisional models of decoherence in, e.g., Joos & Zeh (1985), Kiefer & Joos (1999), and references therein. Furthermore, the conditions for a quantum system to be considered a classical controller of another quantum system support the condition to have determinate values according to EnDQT, since our evidence for measurement-like interactions are based on these situations. [^47] Milburn (2012) provided two examples of interacting quantum systems where one system serves as a classical controller for the other. The conditions necessary for this to occur are that, first, the controller must be open to the environment to establish a pointer basis for the controller coupled with the target system. Second, the dynamics of the controller, as an open system, must ensure that the approach to that pointer basis is much faster than the timescales of the system being controlled. All these conditions support the CDC. So, so far, the relativistic condition seems to be favored. It would be interesting if we find further evidence for or against it.
@@
Before ending this section, I would like to mention two features of this view. First, we can see that EnDQT provides a new interpretation of Born probabilities. They allow us to predict how SDCs evolve. Second, given the above CDCs, how to build models of SDCs? One way is via what I will call the recursive heuristic: as we know, target systems of decoherence models can be environmental systems of other decoherence models. So, given this heuristic, we should consider that (pragmatic irreversible) decoherence models don’t only model measurement-like interactions of the target system, but also how that target system can constitute an environment that gives rise to further measurement-like interactions. So, like in the above simple example involving $A$, $B$, and $C$ that makes certain assumptions, we can then build models of the behavior of SDCs piecewise. If we had the following SDC, $X\rightarrow Y\rightarrow Z$, we should consider at least a decoherence model where $X$ decoheres $Y$ and another one where $Y$ decoheres $Z$. Note that given the CSC4), SDCs have a certain structure. I have represented a toy example of such a structure via Figs. 1 and 2. So, ideally, given CSC4) we should also analyze the subsystems of these systems. Given the recursive heuristic and the fact that EnDQT doesn’t modify the basic equations of QT, it should be possible to develop more realistic models of the evolution of SDCs. [^48] Note also that given the CDCs and EnDQT perspective on quantum states, we can’t infer directly from a system whose quantum state is in an eigenstate of some observable that it has a determinate value of that observable. For instance, when Alice measures her system and assigns it an eigenstate of some observable, she can’t infer that the target system of Bob is in an eigenstate of some observable since Bob might not have interacted with the local SDC. Also, we might assign out of convenience and idealization to a system an eigenstate of some observable, where the latter doesn’t belong to an SDC, but that doesn’t imply that it has a determinate value of that observable. This just implies that if it interacted with an SDC, it could have a determinate value with $100\%$ of probability. Moreover, even upon a measurement of a local system ”in an eigenstate of some observable,” the system shortly after evolves into a superposition (modulo quantum Zeno-like measurements, which increase the probability of the system being found in the same quantum state in repeated measurements). That is also why I have been using decoherence to model measurement-like interactions in general. Furthermore, we can’t always infer from a system that it’s not in an eigenstate of some observable, that it hasn’t a determinate value of such observable. For instance, entangled states that arise in decoherence don’t correspond to eigenstates of some observable. These inferences based on EnDQT are in tension with both directions of the famous Eigenstate-Eigenvalue link because this link neglects the SDCs to make such inferences: A system $S$ has a determinate value $q$ of an observable $O$ if and only if the quantum state of $S$ is in an eigenstate of $O$ with an eigenvalue $q$.
@@
As we can see, EnDQT has the great benefit of achieving UT), providing criteria for absolute determinate values to arise in a single world without modifying the fundamental equations of QT. It only uses decoherence to assign determinate values to a system and the SDCs, whose description appeals to such equations. Furthermore, arbitrary systems can, in principle, be placed in a superposition for an arbitrary time duration concerning any observable as long as they don’t interact with members of an SDC. Also, we have seen that EnDQT provides a series of benefits and predictions.
@@ 3 Why is EnDQT local?
In this section, I will argue that EnDQT achieves LC) and NR) by showing how it provides a local common cause explanation of quantum/Bell-type correlations without adopting non-local/action at a distance, relational, or superdeterministic/retrocausal strategies. In the EPR-Bell scenario, space-like separated Alice and Bob share a pair of quantum systems in an entangled state, and randomly perform measurements on those systems.
First, like in standard QT, the Hamiltonians of interaction, representing the interactions between the agents and their systems, (should) represent local interactions. Second, EnDQT doesn’t modify the equations of QT, and so, in principle, it can be rendered Lorentz invariant, and thus, it can be rendered compatible with relativity and local in this sense. [^49] Bracketing issues with relativistic symmetries that may arise if we aim for a quantum theory of gravity, but this goes beyond the desideratum LC). I will assume this here. Granting that EnDQT achieves these two senses of local, I will also argue that EnDQT deals with the EPR-Bell scenarios without violating relativistic causality, i.e., without forcing us to assume that the causes of correlations travel at a speed faster than the speed of light (the causes of events are always in their past lightcone), and it does that also by providing a local common cause explanation of quantum correlations. Let’s see how. [^50] Future work will enter into more details about this strategy (Pipa, forthcoming-b).
A widely influential version of Bell’s theorem [^51] Another widely influential version is considered to rule out the existence of local deterministic hidden variables (Bell, 1964). On the other hand, the version presented here is often considered to rule out indeterministic hidden variables. Since EnDQT is an indeterministic theory, this version is more relevant. involves, together mainly with the no-superdeterminism assumption, [^52] This assumption states that any events on a space-like hypersurface SH are uncorrelated with any set of interventions subsequent to SH. This theorem also assumes that there is a joint probability distribution for the outcomes of Alice and Bob. the factorizability condition,
@@
The variables $A$, $B$, $\boldsymbol{\Lambda}$, $X$, and $Y$ concern events embedded in a Minkowski spacetime. A and B represent the different measurement results of Alice and Bob, $X$ and $Y$ are the different possible choices of measurement settings for Alice and Bob. $\boldsymbol{\Lambda}$ represents some set of (classical) "hidden" variables in the past lightcone of $A$ and $B$ (see also Figure 3), representing the common causes of the correlations between $X$ and $Y$.
This condition is seen as a consequence of two assumptions: [^53] Bell (1976, 1995, 2004). See also, e.g., Myrvold et al. (2021) and references therein.
  • -The causes of an event are in its past lightcone,
  • -The (what I will call) classical Reichenbach’s Common Cause Principle (CRCCP).
Briefly, the CRCCP states that if events $A$ and $B$ are correlated, then either $A$ causes $B$, or $B$ causes $A$, or both $A$ and $B$ have common causes $\boldsymbol{\Lambda}$, where conditioning on $\boldsymbol{\Lambda}$, $A$ and $B$ are decorrelated, i.e., $P(A,B\mid\boldsymbol{\Lambda})=P(A\mid\boldsymbol{\Lambda})P(B\mid\boldsymbol{\Lambda})$. However, it’s unclear whether we should accept that these probabilistic relations given by the CRCCP should, in general, represent a causal structure involving quantum systems, given the exotic features of the latter. The CRCCP can be seen as a consequence of the Classical Markov Condition (CMC), assumed by classical causal models (CCMs). [^54] I will not derive it here, but see Hitchcock & Rédei (2021).
The CMC connects the causal structure provided by some theory represented by a DAG with probabilistic statements. The CMC is the following,
let’s assume we have a DAG G, representing a causal structure over the variables $\boldsymbol{V}=$$\left\{X_{1},\ldots,X_{n}\right\}$. A joint probability distribution $P\left(X_{1},\ldots,X_{n}\right)$ is classical Markov with respect to $\mathrm{G}$ if and only if it satisfies the following condition: for all distinct variables in $\boldsymbol{V},P$ over these variables factorizes as $P\left(X_{1},\ldots,X_{n}\right)=\prod_{j}P\left(X_{j}\mid Pa\left(X_{j}\right)\right)$, where $Pa\left(X_{j}\right)$ are the "parent nodes" of $X_{j}$, i.e., the nodes whose arrows that start from these nodes point to $X_{j}$.
@@
The acceptability of the CRCCP can be supported by the empirical success of the application of the CMC via CCMs (e.g., Pearl, 2009). [^55] There is also a way of deriving the factorizability condition, as well as the no-superdeterminism condition directly from CCMs and the CMC. See Khanna et al. (2023). EnDQT responds to Bell’s theorem by rejecting that the CMC can be applied in general to accurately represent causal relations between quantum systems, and hence it rejects the applicability of the CRCCP and the factorizability condition to make such accurate representation. [^56] So, note that EnDQT also rejects outcome independence and parameter independence that can be used to derive factorizability condition (Jarrett, 1984) by rejecting their applicability to represent causal relations between quantum systems.
There are at least two complementary ways of justifying the rejection of the CMC. One way is by looking at a precise justification of the CMC involving structural equations: they involve relationships between endogenous variables $V_{j}$ (i.e., variables whose values are determined by other variables in the model) that depend on their endogenous parent variables $\operatorname{Pa}\left(V_{j}\right)$ plus exogenous variables $U_{j}$ (i.e., variables whose values are determined from outside the model) establishing a directed deterministic relationship $V_{j}=f\left(Pa\left(V_{j}\right),U_{j}\right)$. Pearl and Verma (1995) proved that if we have a DAG G’ representing the causal structure on $V_{j}$, the probability distribution $P\left(V_{j}\right)$ that results from the marginalization of the noise sources, if $U_{i}$ are probabilistically independent in P, will respect the $\mathrm{CMC}$ concerning $\mathrm{G}$ [^57] See, e.g., Pearl (2009), Pearl & Verma (1995), and Hitchcock (2022).
The above justification invokes features rejected by EnDQT. First, the systems assumed by these equations have, in general, determinate values in the sense of not having indeterminate ones represented by QT, such as via subsystems of entangled states. Second, the origin of the probabilities of the CMC is in the ignorance about some underlying determinate values. Furthermore, note that these systems with quantum indeterminate values, according to EnDQT, that travel to each wing don’t even have a probabilistic model independently of the measurements of Alice or Bob. So, we can’t have a probability over the common causes independently of their interactions, as it’s assumed by this proof. Third, the above causal relations between systems aren’t described by QT. More precisely, they don’t involve unitary evolutions, decoherence, and quantum indeterministic processes.
We could assign a determinate value to the whole state $|\Psi\rangle$ of the entangled systems that would correspond to the eigenvalues of the observable that this state is an eigenstate of. However, Alice and Bob rather act on the subsystems of these systems. So, we should consider that it is not the whole state $|\Psi\rangle$ that determines the outcomes, but its subsystems. Each subsystem of this entangled state influences locally the outcomes of Alice and Bob, and there is no way to assign a determinate value to each subsystem.
There are at least two possible objections to this justification for the rejection of the CMC. First, there could be some other justification for the CMC that doesn’t assume determinate values, but quantum indeterminate ones, although it’s difficult to see which one that would be. Second, this justification makes it unclear whether a causal explanation of quantum correlations can be provided, and this deficiency could press us to reject other assumptions instead of the CRCCP.

### v5

@@ 2.2 Two hypotheses
One might also object that the inflationary paradigm has its problems, which places EnDQT in problematic foundations. [^67] See, e.g., [51, 30] and references therein. However, note that the SDCs-starting-hypothesis is a placeholder for current and future physics. Furthermore, whatever theory substitutes inflation, it should deal with the problems it pertains to solving. [^68] Note that EnDQT could, in principle, incorporate a cyclic cosmological picture. The early stages refer to the early stages of this universe. To solve those problems, it is plausible to consider that some features shared with the initiators will likely arise. This is because it is plausible to expect that in a future theory, it is likely that we will also have a rapid expansion of the universe in the early stages that doesn’t occur anymore. Such rapid expansion is likely due to some set of entities E with a considerable influence that don’t manifest themselves anymore or so much at least in this way. Furthermore, the SDCs-starting hypothesis can also be supplemented with another hypothesis if it turns out that there are phenomena that need to be explained via widespread initiators. So, EnDQT, conceived more broadly, is a view whose correctness doesn’t just depend on inflation. [^69] It is thus conceivable that there are other alternative mechanisms beyond inflation.
In other words, the SDCs-decoherence hypothesis states that the SDCs in our world are such that the open environments involved in irreversible decoherence pragmatic models will give rise to many interactions between the target systems and systems that have the DC, allowing us to reliably approximate fundamental decoherence models using these pragmatic irreversible decoherence models. From now on, I will just assume that processes represented by pragmatic irreversible decoherence models occur in open environments. Under the plausible assumption that irreversible pragmatic decoherence models represent how the classical domain arises from quantum systems and that this classical domain is widespread, notice that the SDCs-decoherence hypothesis hypothesizes that systems tend to develop interactions that obey CDC1)-CDC4).
The fact that we have, on the one hand, fundamental decoherence models and, on the other hand, pragmatic decoherence models shouldn’t be something objectionable if we accept other QTs because it isn’t, in a sense, a feature exclusive to EnDQT. In the MWI and Bohmian mechanics, we also have, on one hand, the universal wavefunction and, on the other hand, the effective wavefunction. [^70] See, e.g., [44]. Local observers tend to use the effective wavefunction locally and pragmatically, but its predictive success is grounded on the universal wavefunction. Analogously, we use pragmatic decoherence models, but their success is grounded in the phenomena represented by fundamental decoherence models.
One may also worry that there seems to be no guarantee that CDC1)-CDC4) are obeyed by enough systems so that SDCs tend to be formed. As I have said, under the plausible assumption that irreversible pragmatic decoherence models represent phenomena that are widespread, notice that the SDCs-decoherence hypothesis hypothesizes that this is the case. This is plausible because irreversible pragmatic decoherence models represent classical phenomena, and we have evidence that this process is widespread at certain scales. More concretely, we have evidence for irreversible decoherence being widespread because of our evidence concerning entanglement with large macroscopic objects being something widespread. [^71] In other words, systems measured by macroscopic devices can be represented unitarily via this process.
However, one may need more detailed evidence. Note that it is generically true in universes like ours that decoherence qua entanglement tends to increase over time, except under special conditions. Therefore, since CDCs are based on decoherence, the above worry shouldn’t be a problem because we have grounds to believe that more systems will follow CDC1)-CDC4) as time evolves. [^72] Furthermore, there is a related plausible hypothesis that can be regarded at least in part as a consequence of a particular specification of the SDCs-starting hypothesis and the SDCs-decoherence hypothesis, which I think also defeats this skepticism. I will outline this hypothesis here and the argument that appeals to it and leave its details for future work. It has been argued by many, such as one of the pioneers of the decoherence program ([107], see also [4] and references therein), that decoherence gives rise to its own arrow of time. Roughly, due to the irreversibility of decohering interactions, quantum states of systems in the universe tend to be more mixed in the future than in the past. Although EnDQT doesn’t reify quantum states, given the SDCs-decoherence hypothesis, it interprets irreversible pragmatic decoherence as standing for a particular physical phenomenon, and thus, an analogous process should occur for EnDQT. Since CDCs are based on decoherence, we can instead consider that there is a tendency for more systems to follow the CDCs in the future than in the past, forming more SDCs in the future than in the past. More concretely, we can consider an SDCs-starting hypothesis that establishes that there were no SDCs and just initiators in the past, unentangled with other systems (as it is standardly assumed by decoherence models). Then, as the systems evolve under physically plausible Hamiltonians, there will be a tendency for more systems to interact, get entangled, give rise to decoherence, and form interactions that obey the CDCs, giving rise to SDCs. If we adopt the inflationary-starting hypothesis, we may end up considering that this progressively increasing number of systems that will tend to belong to SDCs at some point is accompanying the beginning and evolution of the universe. Thus, what I will call SDCs-time’s arrow hypothesis should be seen as a consequence of at least a specification of the SDCs-starting hypothesis and the SDCs-decoherence hypothesis. As we can see, this argument doesn’t assume anything about decoherence being widespread. Thus, if we also endorse this hypothesis, the above worry shouldn’t be a problem, and in principle, we have grounds to believe that more systems will follow CDC1)-CDC4) as time evolves.
One may also object to the following: suppose agent $M$ interacts with $S$ such that $S$ acquires a determinate value for some binary observable. Then agent $M$ sends this value to $M^{*}$ using, say, the z-spin of a single particle as a signal. There is a period during which the signal particle isn’t interacting with anything, and thus, that system $S$ doesn’t belong to an SDC. Therefore, it has an indeterminate value of the observable whose value represents the information $M$ wants to communicate to $M^{*}$. Given this, it is hard to conceive how $S$ can carry reliable information about the result that $M$ got without itself having a determinate value.
As a reply, note that a system $S$ can carry reliable information if the complex amplitude $\alpha$ that is associated with the quantum state that concerns the information that we want to transmit is such that $|\alpha|^{2}>>1$. When we approach EnDQT, we need to deidealize quantum mechanics. Systems, after measurements upon their self-Hamiltonian, tend to evolve out of the quantum state that was measured, where this quantum state could encode information that aims to be sent somewhere else. As mentioned before, the assumption that systems have indeterminate values of any observable by default except during certain interactions comes from the perspective that the so-called links (like the EEL) that connect state assignments to assignments of determinate values are idealizations that never occur in practice.
@@
I have mentioned above some predictions that EnDQT provides, such as the dissolution of SDCs and those that are a consequence of the SDCs-inflationary hypothesis. Another prediction is the following: as we have seen with the example above, adopting the CDCs generates constraints on how SDCs are formed and new predictions. Decoherence timescales roughly serve as an indicator for the timescale it takes for environments of a system to decohere that system on average, where that system ends up having specific determinate values (that are observed in the lab). Given the SDCs-decoherence hypothesis, the CDCs predict that the decoherence timescale that we empirically observe of a kind of system $Z$ by a kind of system $Y$ should be superior or of the same order as the decoherence timescale of $Y$ by a kind of system $X$, where $Y$ is typically decohered by $X$ before $Y$ decoheres $Z$, and where the interaction between $X$ and $Y$ starts first. Otherwise, contrary to what is assumed by the CDCs, we can have situations where $Z$ will have a determinate value first (due to $Y$), then $Y$ will have a determinate value due to $X$. Since the decoherence timescales can be empirically determined, a further analysis of the current empirically determined decoherence timescales is needed to see if they agree with the predictions of the CDCs.
@@
The predictions of the CDCs are supported in the case that $Y$ is a macrosystem (e.g., measurement devices), and $Z$ is a microsystem. This is because macroscopic systems have decoherence timescales much shorter than the microscopic systems that they can decohere. [^73] The cross-section for larger systems is larger than the one for smaller systems. Moreover, the decoherence rate of a quantum system, which is the inverse of the decoherence timescale, is proportional to their cross-section, as well as the flux of systems of the environment. See the collisional models of decoherence in, e.g., [53, 55, 89], and references therein. Furthermore, the conditions for a quantum system to be considered a classical controller of another quantum system support the CDCs, since our evidence for measurement-like interactions are based on these situations.747474 [68] provided two examples of interacting quantum systems where one system serves as a classical controller for the other. The conditions necessary for this to occur are as follows: First, the controller must be open to the environment to establish a pointer basis for the controller coupled with the target system. Second, the dynamics of the controller, as an open system, must ensure that the approach to that pointer basis is much faster than the timescales of the system being controlled. All these conditions support the CDCs. So, so far, the CDCs seem to be favored. It would be interesting if we find further evidence for or against them.
@@
Before ending this article, I would like to mention two additional features of this view. First, EnDQT provides a new interpretation of Born probabilities, viewing them as predictors of how SDCs evolve upon specific interactions. [^75] If one is committed to objective probabilities qua chances, they track those chances that arise under these interactions Second, given the above CDCs, we can adopt different strategies to build models of SDCs. One strategy is what I will call the recursive heuristic: as we know, target systems of decoherence models can be environmental systems of other decoherence models. So, given this heuristic, we should consider that (pragmatic irreversible) decoherence models don’t only model measurement-like interactions of the target system, but also how that target system can constitute an environment that gives rise to further measurement-like interactions. So, like in the above simple example involving $A$, $B$, and $C$ that makes certain assumptions, we can then build models of the behavior of SDCs piecewise. If we had the following SDC, $X\rightarrow Y\rightarrow Z$, we should consider at least a decoherence model where $X$ decoheres $Y$ and another one where $Y$ decoheres $Z$ to describe this process. Note that given the CDC4), SDCs have a certain structure. I have represented a toy example of such a structure in Figs. 1 and 2. In Appendix A, I show in more detail how this can be done. Given the recursive heuristic and the fact that EnDQT doesn’t modify the basic equations of QT, it should be possible to develop more realistic models of the evolution of SDCs. [^76] Note also that given the CDCs and EnDQT perspective on quantum states, we cannot infer directly from a system whose quantum state is in an eigenstate of some observable that it has a determinate value of that observable. For instance, when Alice measures her system and assigns it an eigenstate of some observable, she cannot infer that the target system of Bob is in an eigenstate of some observable since Bob might not have interacted with the local SDC. Also, as I have mentioned above, we might assign out of convenience and idealization to a system an eigenstate of some observable, where the latter doesn’t belong to an SDC, but that doesn’t imply that it has a determinate value of that observable. This just implies that if it interacted with an SDC, it could have a determinate value with $100\%$ of probability. Moreover, even upon a measurement of a local system ´´in an eigenstate of some observable,” the system shortly after evolves into a superposition (modulo quantum Zeno-like measurements, which increase the probability of the system being found in the same quantum state in repeated measurements). That is also why I have been using decoherence to model measurement-like interactions in general. Furthermore, we cannot always infer from a system that it is not in an eigenstate of some observable, that it hasn’t a determinate value of such observable. For instance, entangled states that arise in decoherence don’t correspond to eigenstates of some observable. These inferences based on EnDQT are in tension with both directions of the Eigenstate-Eigenvalue link because this link neglects the SDCs to make such inferences.
@@
Given the CDC1)-CDC4), and the above two hypotheses, EnDQT has the important benefit of achieving UT), providing criteria for absolute determinate values to arise in a single world without modifying the fundamental equations of QT. It only uses decoherence to assign determinate values to a system and the SDCs, whose description appeals to such equations. Furthermore, arbitrary systems can, in principle, be placed in a superposition for an arbitrary duration concerning any observable as long as they don’t interact with members of an SDC. Of course, given the SDCs-decoherence hypothesis, in principle, doesn’t mean in practice. Our pragmatic models of decoherence tell us that it’s very difficult to place large macroscopic systems in a superposition. Also, we have seen that EnDQT provides a series of other benefits and predictions, making it a QT worth taking seriously.
@@ 3 Why is EnDQT local?
In this section, I will argue that EnDQT achieves LC) and NR) by showing how it deals with Bell’s theorem and provides a local common cause explanation of quantum/Bell-type correlations without adopting non-local/action-at-a-distance, relational, or superdeterministic/retrocausal strategies. In the EPR-Bell scenario, space-like separated Alice and Bob share a pair of quantum systems in an entangled state and randomly perform measurements on those systems.
First, like in standard QT, the Hamiltonians of interaction, representing the interactions between the agents and their systems, should represent local interactions. Second, EnDQT doesn’t modify the equations of QT, and so, in principle, its laws can be rendered Lorentz covariant and generally covariant, and thus, it can be rendered compatible with relativity and be local in this sense. More precisely, if QT allows laws to be Lorentz and generally covariant (which seems to be the case at least given Quantum Field Theory in flat and curved spacetime), [^77] See, e.g., [94]. then since EnDQT doesn’t modify standard QT, it should uphold these relativistic symmetries. [^78] More on this in Appendix A. I am bracketing issues with relativistic symmetries that may arise if we aim for a quantum theory of gravity, but this goes beyond the desideratum LC). I will assume this here; however, future work will show this explicitly. Assuming that EnDQT achieves these two senses of locality, I will further argue that EnDQT addresses the EPR-Bell scenarios without violating relativistic causality, in the sense that it doesn’t require us to assume that the causes of the events involved in those correlations are not within their past lightcone, and without invoking superdeterministic or retrocausal explanations. Furthermore, it provides a local common cause explanation of quantum correlations. Let’s see how. I will focus first on the 1976 widely influential version of Bell’s theorem because it is generally considered applicable to indeterministic theories. Since EnDQT is an indeterministic theory, this version is more relevant. [^79] Another widely influential version is considered to rule out the existence of local deterministic hidden variables [9]. I will go over this version further below. This version assumes the so-called statistical independence or no-superdeterminism assumption. This assumption states that any events on a space-like hypersurface are uncorrelated with any set of interventions subsequent to this hypersurface. It also assumes that there are single observed outcomes and not, for example, multiple outcomes that correspond to multiple worlds or perspectives. [^80] Another way of putting this assumption is that there is a joint probability distribution involving the outcomes of Alice and Bob. See the absolutness of observed events assumption further below. This is the assumption denied, for example, by the Many-Worlds Interpretation. The assumption more relevant for EnDQT is the factorizability condition. According to this condition,
@@
The variables $A$, $B$, $\boldsymbol{\Lambda}$, $X$, and $Y$ represent events embedded in a Minkowski spacetime. $A$ and $B$ represent the different measurement results of Alice and Bob, while $X$ and $Y$ are the different possible choices of measurement settings for Alice and Bob. $\boldsymbol{\Lambda}$ represents some set of (classical) “hidden” variables in the past lightcone of $A$ and $B$ (see also Figure 3), representing the common causes of the correlations between $X$ and $Y$.
This condition arises as a consequence of two assumptions:818181[10, 11]. See also, e.g., [72] and references therein.
  • The causes of events are always in their past lightcone, [^82] By always, we can also mean in any relativistic reference frame.
  • The Classical Reichenbach Common Cause Principle (CRCCP).
I will be concerned here with a version of the CRCCP that is expressed in terms of variables whose different values represent different values of certain quantities or physical features. This is because this is the most appropriate notion to represent features of at least classical physical systems (i.e., systems represented via classical mechanics). [^83] There are other versions that use propositions and their negation to state this principle (i.e., “events”). See, .e.g., [49]. Briefly, this version of the CRCCP states that if variables $A$ and $B$ are correlated, then either $A$ causes $B$, or $B$ causes $A$, or both $A$ and $B$ have common cause variable $\boldsymbol{\Lambda}$, where conditioning on $\boldsymbol{\Lambda}$, $A$ and $B$ are decorrelated, i.e., $P(A,B\mid\boldsymbol{\Lambda})=P(A\mid\boldsymbol{\Lambda})P(B\mid\boldsymbol{\Lambda})$. However, it is unclear whether we should accept these probabilistic relations given by the CRCCP as generally representing a causal structure involving quantum systems, given the exotic features of quantum systems. The above version of the CRCCP can be derived from the Classical Markov Condition (CMC), assumed by the so-called Classical Causal Models (CCMs). [^84] See Appendix E for this derivation. See also [49]. Note that the version of the CRCCP mentioned in the previous footnote cannot be derived from the CMC.
The CMC connects the causal structure provided by some theory, and which is represented by a Directed Acyclic Graphs (DAGs), i.e., a directed graph with no cycles, with probabilistic statements. Note that the sense “causation” will be understood merely in terms of influence between quantum system, represented via QT (more on this below). The CMC says the following,
Let’s assume we have a DAG G, representing a causal structure over the variables $\boldsymbol{V}=$$\left\{X_{1},\ldots,X_{n}\right\}$. A joint probability distribution $P\left(X_{1},\ldots,X_{n}\right)$ is classical Markov with respect to $\mathrm{G}$ if and only if it satisfies the following condition: for all distinct variables in $\boldsymbol{V},P$ over these variables factorizes as $P\left(X_{1},\ldots,X_{n}\right)=\prod_{j}P\left(X_{j}\mid Pa\left(X_{j}\right)\right)$, where $Pa\left(X_{j}\right)$ are the “parent nodes” of $X_{j}$, i.e., the nodes whose arrows point to $X_{j}$.
@@
Given the widespread empirical success of the application of the CMC via CCMs (e.g., [78]), which can be used to derive the CRCCP, I will consider that the empirical success of the CRCCP in physics is supported by the empirical success of the application of the CMC via CCMs rather than the other way around (e.g., [78]). [^85] There is also a way of deriving the factorizability condition, as well as the no-superdeterminism condition directly from CCMs and the CMC. See [54]. EnDQT responds to Bell’s theorem by rejecting that the CMC, and hence the CCMs, can be applied in general to accurately represent causal relations between quantum systems, and hence, it rejects the applicability of the CRCCP and the factorizability condition to make such an accurate representation. [^86] Therefore, note that EnDQT also rejects outcome independence and parameter independence [52], which can be used to derive the factorizability condition by rejecting their applicability to represent causal relations between quantum systems. This argument is schematized in 4.
Figure 4: Diagram that helps to understand how EnDQT deals with Bell’s 1976 theorem, which is considered to be applicable to indeterministic theories. It also helps in understanding how EnDQT deals with it. It involves the rejection of the applicability of the Causal Markov Condition and Classical Causal Models to, in general, accurately represent causal relations between quantum systems. If a box with the claim $X$ is connected by one or more arrows, means that the one or more boxes with claims $Y$ that point to $X$ jointly lead to the derivation of $X$.
There are at least two complementary arguments one can give to justify the rejection of the CMC, which aren’t necessarily mutually exclusive. One argument looks directly at (to my knowledge) the most precise derivation of the CMC via structural equations to conclude that the CMC and the CCMs are inappropriate to represent causal relations between quantum systems. These equations involve relationships between endogenous variables $V_{j}$ (i.e., variables whose values are determined by other variables in the model) that depend on their endogenous parent variables $\operatorname{Pa}\left(V_{j}\right)$ plus exogenous variables $U_{j}$ (i.e., variables whose values are determined from outside the model) establishing a directed deterministic relationship $V_{j}=f\left(Pa\left(V_{j}\right),U_{j}\right)$. [79] proved that if we have a DAG G’ representing the causal structure on $V_{j}$ (i.e., a “causal DAG”), [^87] This DAG is often called a Bayesian network. the probability distribution $P\left(V_{j}\right)$ that results from the marginalization of the noise sources if $U_{i}$ are probabilistically independent in P, will respect the $\mathrm{CMC}$ concerning $\mathrm{G}$ [^88] See, e.g., [78, 79, 49]. The above derivation assumes features rejected by EnDQT. First, the origin of the probabilities of the CMC is in the ignorance about some underlying determinate values. Furthermore, note that these systems that travel to each wing, which, according to EnDQT, have indeterminate values, don’t even have a probabilistic model independently of the measurements of Alice or Bob. [^89] Note that we could assign a determinate value to the whole state $|\Psi\rangle$ of the entangled systems that would correspond to the eigenvalues of the observable that this state is an eigenstate of. However, Alice and Bob rather act on the subsystems of these systems. So, we should consider that it is not the whole state $|\Psi\rangle$ that determines the outcomes but its subsystems. Each subsystem of this entangled state influences locally the outcomes of Alice and Bob, and there is no way to assign a determinate value to each subsystem. So, we cannot have a probability over the common causes independently of their interactions, as it is assumed by this proof. Third, the above causal relations between systems aren’t described by QT. More precisely, they don’t involve unitary evolutions, decoherence, and quantum indeterministic processes. Further below, I will show how EnDQT, by allowing indeterminate values, also permits a local explanation of Bell correlations. Given that this precise derivation of the CMC makes assumptions that are rejected by EnDQT, I, therefore, conclude that the CMC and CCMs are inadequate to represent causal relations between quantum systems according to EnDQT.